OpenStudy (anonymous):

Evaluate x^-2 for x = -3. I don't know how to solve this?

OpenStudy (tkhunny):

Have you considered making the substitution?

jigglypuff314 (jigglypuff314):

as in plugging in -3 for x into x^-2 so (-3)^-2 note that the negative exponent means that the value flips so like 3^-1 = 1/3 x^-3 = 1/(x^3) 5^-2 = 1/(5^2) = 1/25 so on and so forth

OpenStudy (jdoe0001):

\(\bf a^{-{\color{red} n}} = \cfrac{1}{a^{\color{red} n}} \\ \quad \\ \quad \\ {\color{red}{ x}}^{-2}\qquad {\color{red}{ x}}=-3\implies ({\color{red}{ -3}})^{-2}\implies \cfrac{1}{({\color{red}{ -3}})^{2}}\implies ? \)